OpenStudy (callisto):

What does \[e^{i\pi}\] mean? What is its significance? Why we have to define such a number?

OpenStudy (unklerhaukus):

negative

OpenStudy (anonymous):

one

OpenStudy (anonymous):

i still don't know why people love it so much. maybe i can't see what they see.

OpenStudy (anonymous):

it's basically the relationship of the three constants.

OpenStudy (anonymous):

e, iota and pi

OpenStudy (earthcitizen):

Euler's formula

OpenStudy (lalaly):

\[e^{i \theta}=\cos \theta +i \sin \theta\]

OpenStudy (anonymous):

And the you have it as your pic, without knowing what does it mean?

OpenStudy (unklerhaukus):

that is no iota, iota does not have a dot

OpenStudy (callisto):

@Ishaan94 I just think it looked interesting, with 3 special numbers in one term. I knew nothing about it when I uploaded it...

OpenStudy (anonymous):

Use @lalaly formula and instead of Theta, substitute pi you will get -1

OpenStudy (anonymous):

hmm \(\sqrt{-1}\) isn't called iota?

OpenStudy (anonymous):

but that becomes a complex number

OpenStudy (unklerhaukus):

call it i

OpenStudy (unklerhaukus):

as in eye, for imaginary

OpenStudy (anonymous):

ohh yeah, indian education system sucks or teachers or maybe just me. my life's been a lie. it's not iota! but they made me think it is.

OpenStudy (unklerhaukus):

well im sorry to here that but i kinda like the fact the were all Greek for a moment there

OpenStudy (anonymous):

iota = root of -1

OpenStudy (anonymous):

@Ishaan94 , iota is the greek pronunciation of i

OpenStudy (anonymous):

\(\iota\) this 'iota' shivam.

OpenStudy (callisto):

iota is a greek letter(with no dot)... sqrt -1 is i.... as I've learnt..

OpenStudy (anonymous):

is*

OpenStudy (unklerhaukus):

this is iota and there is no dot on it \[ \iota \]

OpenStudy (anonymous):

lol really?

OpenStudy (anonymous):

LOLOL The teachers in India have to be tested for their knowledge because they are passing wrong knowledge to us too :(

OpenStudy (anonymous):

haha

OpenStudy (callisto):

What is its significance?

OpenStudy (anonymous):

See If you multiply a complex number with e^i(Theta) Then the line rotates about angle theta to anticlockwise direction. Diagam below'  |z_2|=|z_1|

OpenStudy (anonymous):

Therefore z_2 = z_1 * e^(i(theta))

OpenStudy (anonymous):

it means it's too difficult to comprehend

OpenStudy (anonymous):

the limit of human intellect

OpenStudy (callisto):

I can see that it would be something interesting...

OpenStudy (anonymous):

it means that there's a creator of the universe

OpenStudy (anonymous):

Have a look at complex numbers. If you master it , it would make your life and maths and physics problem solving a hell lot easier. :)

OpenStudy (unklerhaukus):

i something to do with phase

OpenStudy (anonymous):

\( i \quad \iota \quad j\) are some usual notation of the imaginary units. I would never blame my teachers over some notation.

OpenStudy (anonymous):

http://isomorphismes.tumblr.com/post/754631763/eipi

OpenStudy (anonymous):

@FoolForMath , Ok that brings some clarity to me . Thanks :D @estudier , an awesome link :D @everyone, I would say that Euler's law is one of the most important and most useful law I have come across in maths. it makes your life lot easier.

OpenStudy (callisto):

The link helps a lot. Thanks! :)

OpenStudy (anonymous):

It's just beautiful, don't ask me why .. it's just is.

OpenStudy (anonymous):

Can I ask How? lol hmm -sigh-

OpenStudy (callisto):

e, pi => irrational i -> not real when they come to together and form e^( i pi) = -1 -1 is rational and real irrational + not real -> rational and real => beauty :)